OpenStudy (anonymous):

If tan(x+y)=x then dy/dx= Answer: cos^2(x+y)-1 Please show how to get to this answer.

hartnn (hartnn):

do u know chain rule ?

OpenStudy (anonymous):

yes i do

hartnn (hartnn):

so where are you stuck ? could you start ?

OpenStudy (anonymous):

I got this far...

OpenStudy (anonymous):

\[\sec ^{2}(x+y)(1+\frac{ dy }{ dx })\] \[\sec ^{2}(x+y)(1+\frac{ dy }{ dx })-1\] and i was kinda confused where to go from there

hartnn (hartnn):

our main aim is to isolate dy/dx distribute sec^2 (x+y) inside brackets

OpenStudy (anonymous):

ok so i have... \[[\sec^2x+\sec^2y](1+\frac{ dy }{ dx })-1=0\]

OpenStudy (anonymous):

should i put the 1 back on the other side and divide by the [sec^2...]?

hartnn (hartnn):

that can't be done... sec(A+B) does not = secA +secB i was talking about this \(\sec^2 (x+y)(1+y') = \sec^2(x+y) + y' \sec^2 (x+y)\)

OpenStudy (anonymous):

oh ok. i got it now. thank you :)

hartnn (hartnn):

could you find dy/dx properly ? and welcome ^_^

OpenStudy (anonymous):

yes. i got it :)